OpenStudy (anonymous):

I have 200 dollars to split between 3 teams. Second place gets twice as much as third place and first gets 15 dollars more than second place. Can someone please help me with the formula?

OpenStudy (anonymous):

let x be the amount that the third place got. then 2nd place get 2x and 1st place get 2x+15. sums all of them equal to 200

OpenStudy (anonymous):

wouldn't take 200 /3+66.66 the at 15 to the first team would have 81.66

OpenStudy (anonymous):

200/3=66.66

OpenStudy (anonymous):

thrid would have 66.66 and two ther teams would have 15 add two this 66.66

OpenStudy (anonymous):

hopefuly maybe help you out

OpenStudy (anonymous):

the formula: x+2x+(2x+15)=200 [where x is the amount 3rd place get] 5x+15=200 x=37 therefore, 2nd place wil get, 2x=2*37=74 and 1st place gets 2x+15=74+15=89

OpenStudy (anonymous):

Thanks anfar

OpenStudy (anonymous):

no problem.